# On the non-Markovian quantum control dynamics

### v2

@@ II-B4 Example
Take the parameter settings in Section II-B3) as an example. When $\Omega=\omega_{a}$ and $\frac{R}{P}-\left(\frac{Q}{2P}\right)^{2}=0$, $F_{I}(t)\equiv 0$, $Q=-\gamma$, $R=\frac{\gamma^{2}}{4P}$, $F_{R}(t)=\frac{\gamma}{2P}-\frac{\gamma}{\gamma Pt+2P}$, then $PF_{R}(t)-\gamma<0$ for arbitrary $t$ and $\textbf{J}\left(X_{F}(t)\right)$ is definitely negative. If we take $\mathbf{M}=\mathbf{I}$ for simplification in Eq. (32), then $\dot{V}(X_{F})<0$ and Eq. (24) is globally stable at this parameter setting.
Definition 3 (see [66]).
A set $S$ is said to be invariant if each solution starting in $S$ remains in $S$ for all $t$.
Generalized from the stability property in Ref. [66], we can derive the following proposition.
@@
For the set $\Omega_{\alpha}={X_{F}\in\textbf{R}^{2}:V\leq\alpha_{V}}$ is bounded for $\alpha_{V}>0$, when $02\sum_{n}|\kappa_{n}|^{2}\alpha_{V}+\gamma\chi-2\gamma<$ 0 and $\mathbf{M}=\mathbf{I}$, then a local invariant set contained in the region $\dot{V}=0$ and the scale of the invariant set is determined by the parameters as $\alpha_{V}\leq\frac{\gamma(2-\chi)}{2\sum_{n}|\kappa_{n}|^{2}}$ with $20<\chi<$ 2.
@@
Assume initially $X_{F}\in\Omega_{\alpha}$ and $\Gamma_{F}$ represents an invariant set of $X_{F}$ satisfying that $\dot{V}(X_{F})=0$. When $02\sum_{n}|\kappa_{n}|^{2}\alpha_{V}+\gamma\chi-2\gamma<$ 0, then $\dot{V}<0$ can be satisfied according to Eqs. (32), (24a) with $\Omega=\omega_{a}$, and $V(x)\geq 0$ by its definition. Then there exists a positive value $v_{l}$ that $\lim_{t\rightarrow\infty}V(X_{F})=v_{l}\geq 0$ and $v_{l}<\alpha_{v}$. Thus $V(X_{F})$ is always within the invariant set. ∎
@@
Proposition 2 means that larger $\gamma$ or smaller $\sum_{n}|\kappa_{n}|^{2}$ can induce a larger invariant set with stronger Markovian property. This is why the Markovian approximation can be applied in the circumstance that the coupling between the quantum system and the environment is weak.
Proposition 3 ([63]).
Suppose for each $u\geq 0$, there exists a positive Lyapunov function $V_{u}$ for the system with the control $u$ and
$\dot{V}_{u}(t,X_{f})\leq 0.$
When the control satisfies $\|u\|\leq a_{u}$, then the system is BIBO stable.
@@
The system in Eq. (24) is BIBO stable when the initial condition $\left\|F(0)\right\|$ is bounded and the system starts in the invariant set $\Omega_{\alpha}={X_{F}\in\textbf{R}^{2}:V\leq\alpha_{V}},$ satisfying $02\sum_{n}|\kappa_{n}|^{2}\alpha_{V}+\gamma\chi-2\gamma<$ 0.
@@
When the condition is satisfied, then $\dot{V}<0$ according to Eq. (32), then the system in Eq. (24) is BIBO based on Proposition 3. ∎
@@
When the system is BIBO, the Jacobian in Eq. (28) is also bounded, then the Lyapunov function can also be bounded according to Ref. [67].
@@
When $\dot{V}<0$, then $\lim_{t\rightarrow\infty}F_{R}(t)$ and $\lim_{t\rightarrow\infty}F_{I}(t)$ exist when $\left|\Omega-\omega_{a}\right|\leq\epsilon$.
@@
When $\dot{V}<0$ and $V(t)>0$, then $\lim_{t\rightarrow\infty}V(t)$ exists and $\lim_{t\rightarrow\infty}\dot{V}(t)=0$. Then $\lim_{t\rightarrow\infty}F_{R}(t)$ and $\lim_{t\rightarrow\infty}F_{I}(t)$ can be solved by the first line of Eq. (32). ∎
Then we can derive the following proposition based on the results above and the following assumption [68].
@@
Assume that the Lyapunov function in Eq. (32) satisfies $\gamma_{1}\left|X_{F}(t)\right|^{2}\leq V\left(X_{F}(t)\right)\leq\gamma_{2}\left|X_{F}(t)\right|^{2}$.
Proposition 5.
If the system with $\Omega=\omega_{a}$ is BIBO stable and $\left|\Omega-\omega_{a}\right|\leq\epsilon$, then the system will finally converge to the Markovian case as $t\rightarrow\infty$.
@@
The Lyapunov function in Eq. (32) is independent from $\left(\Omega-\omega_{a}\right)$. When the system is BIBO stable and $\left|\Omega-\omega_{a}\right|\leq\epsilon$, then $F_{R}(t)$ and $F_{I}(t)$ are solvable with steady values according to Lemma 1 and Eq. (28). ∎
The real and imaginary parts of $F(t)$ are compared in Fig. 2(a), where we take $\chi=0.5$ and $\gamma=1$. For the solid lines in (a), $P=1$, $\Omega=\omega_{a}=50$, while for the dashed lines, $P=1$, $\Omega=50$ and $\omega_{a}=45$, which means that there is a detuning between the atom and the environment. Besides, the purple dot line is for the case that $P=0.2$ and $\Omega=\omega_{a}=50$, which illustrates that smaller $P$ can induce faster convergence to Markovian interactions with the environment. The bounded evolution of the real and imaginary parts in (a) are further compared in (b) with the red line for $\gamma=1$, the green line for $\gamma=0.5$, the black line for $\gamma=3$, and the other parameters are the same as the dashed lines in (a). The comparisons in (b) show that larger $\gamma$ is better for the realization of Markovian interactions with the environment, which agrees with Remark 3.
II-C An example to clarify how $F(t)$ influences the atomic dynamics
In this subsection, we take the two-level atom as an example. One initially excited two-level atom is coupled with a cavity, and the atom is also coupled with the environment. The dynamics of the system can be represented by Eq. (9) with $N=2$, and the dynamics is compared as follows.
Figure 3: The Markovian and non-Markovian dynamics of the quantum system with one two-level atom in the cavity.
As shown in Fig. 3, initially the atom is excited with $\left\langle\sigma^{+}_{1}\sigma^{-}_{1}\right\rangle=1$ and the other amplitudes of quantum states in Eq. (9) are zero. We take the atom and cavity parameters as $\omega_{a}=50$, $g_{1}=2$ and $\delta=0$. When the atom is coupled with a non-Markovian environment with the parameters $\Omega=45$, $\kappa_{1}=4$, $\chi=1$ and $\gamma=1$, then the population of the atom’s two different states and the number of photons in the cavity are shown as the solid lines in Fig. 3. Besides, a simplified case that the atom interacts with the Markovian environment is simulated as the dashed lines with $F(t)\equiv 0.0792$, which is the final steady amplitude of $F(t)$ in the non-Moarkovian case. The larger oscillations of the solid lines are induced by the detuning between the two-level atom and the environment valued by $\Omega-\omega_{a}$ as in Eq. (9e), then as a result, $F(t)$ can be complex values but always time-varying in the non-Markovian parameter settings.
II-D Nonlinear dynamics with uncertain control input
Practically, we usually cannot know the exact value of the environment parameter $\Omega$, thus the component $\Omega-\omega_{a}$ in Eq. (24) can be regarded as an uncertain control input. Then Eq. (25) can be re-written as
where we denote $\textbf{u}=\Delta AX_{F}(t)$ with $\Delta A=\begin{bmatrix}0&\Omega-\omega_{a}\\ \omega_{a}-\Omega&0\end{bmatrix}$, and there always exists $\epsilon>0$ satisfying $\|\Delta A\|\leq\epsilon$. Then Eq. (33) can be regarded as a perturbed nonlinear system.
Assumption 5.
The nonlinear equation (24) with $\Omega=\omega_{a}$ is stable according to the parameter settings in Proposition 1.
According to Eqs. (18)-(22), when $\Omega=\omega_{a}$, $F_{I}(t)\equiv 0$, and we denote $\lim_{t\rightarrow\infty}F_{R}(t)=\bar{F}_{R}$. Define a new state vector $\tilde{X}_{F}(t)=\left[F_{R}(t)-\bar{F}_{R},F_{I}(t)\right]^{T}\triangleq\left[\tilde{F}_{R}(t),F_{I}(t)\right]^{T}$ and $\lim_{t\rightarrow\infty}\tilde{X}_{F}(t)=\left[0,0\right]^{T}\triangleq\tilde{X}_{F}^{*}$ when Assumption 5 is satisfied and $\Omega=\omega_{a}$. Then
$\frac{\mathrm{d}}{\mathrm{d}t}\tilde{F}_{R}(t)=\sum_{n}|\kappa_{n}|^{2}\left[\left(\tilde{F}_{R}(t)+\bar{F}_{R}\right)^{2}-F_{I}^{2}(t)\right]$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\gamma\left(\tilde{F}_{R}(t)+\bar{F}_{R}\right)+\left(\Omega-\omega_{a}\right)F_{I}(t)+\frac{\gamma\chi}{2},$ (34a)
$\frac{\mathrm{d}}{\mathrm{d}t}F_{I}(t)=2\sum_{n}|\kappa_{n}|^{2}\left(\tilde{F}_{R}(t)+\bar{F}_{R}\right)F_{I}(t)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\left(\Omega-\omega_{a}\right)\left(\tilde{F}_{R}(t)+\bar{F}_{R}\right)-\gamma F_{I}(t),$ (34b)
and can be simplified as
$\frac{\mathrm{d}}{\mathrm{d}t}\tilde{X}_{F}(t)=$ $\tilde{\textbf{f}}\left(\tilde{X}_{F}(t)\right)+\begin{bmatrix}\frac{\gamma\chi}{2}+\sum_{n}|\kappa_{n}|^{2}\bar{F}_{R}^{2}-\gamma\bar{F}_{R}\\ \left(\omega_{a}-\Omega\right)\bar{F}_{R}\end{bmatrix}$ (35)
$+\begin{bmatrix}0&\Omega-\omega_{a}\\ \omega_{a}-\Omega&0\end{bmatrix}\tilde{X}_{F}(t),$
where $\tilde{\textbf{f}}\left(\tilde{X}_{F}(t)\right)$ is for the nonlinear component at the RHS of Eq. (34). Obviously, $\tilde{\textbf{f}}\left(\tilde{X}_{F}(t)\right)=0$ when $\tilde{X}_{F}=0$. When $\Omega=\omega_{a}$ and $X_{F}(t)$ converges by Proposition 1, the RHS of Eq. (35) converges to zero.
Proposition 6 (see [69]).
For the nonlinear system
Proposition 7.
When $\frac{R}{P}-\left(\frac{Q}{2P}\right)^{2}\leq 0$, there exists $\epsilon>0$ such that the quantum system approaches a Markovian behavior when $\left|\omega_{a}-\Omega\right|\leq\epsilon$.
Proof.
@@ II-D Nonlinear dynamics with uncertain control input
When $\omega_{a}=\Omega$, by Proposition 1, the quantum dynamics approaches Markovian behavior as $t\rightarrow\infty,$ and the nonlinear dynamics $\frac{\mathrm{d}}{\mathrm{d}t}\tilde{X}_{F}(t)=\tilde{\textbf{f}}\left(\tilde{X}_{F}(t)\right)$ is uniformly and asymptotically stable around its equilibrium $\tilde{X}_{F}=0$. When $\left|\omega_{a}-\Omega\right|\leq\epsilon$ and $\gamma\chi$ is finite, Proposition 6 ensures that $\tilde{F}_{R}(t)$ and $F_{I}(t)$ in Eq. (34) remain stable. Then the quantum dynamics will approach Markovian behavior as $t\rightarrow\infty$. ∎
@@ III Time-varying linear quantum control dynamics
In the following, we only consider the circumstance that $\frac{R}{P}-\left(\frac{Q}{2P}\right)^{2}\leq 0$, thus the interaction between the atom and environment finally converges to a Markovian form. Return to Eq. (12), we divide the following analysis into two parts according to whether there are detunings between the atom and cavity, as clarified in Remark 2.
III-A $\delta=0$, $g_{n}\neq 0$
At this parameter setting, $C_{n}$ and $D_{n}$ in Eqs. (15) and (16) are all time-invariant for arbitrary $n$, while $A_{n}$ is time-varying with $F(t)$ converges to its steady values according to Eqs. (19) and (22). $A(t)$ can be separated into the time-invariant component and time-varying components as
@@ III-A $\delta=0$, $g_{n}\neq 0$
$=$ $\begin{bmatrix}-L_{n}^{(1)}&-ig_{n}&ig_{n}\\ -ig_{n}&-L_{n}^{(2)}&0\\ ig_{n}&0&-L_{n}^{(3)}\end{bmatrix}$
$+\begin{bmatrix}L_{n}^{(1)}-\left(F(t)+F^{*}(t)\right)|\kappa_{n}|^{2}&0&0\\ 0&L_{n}^{(2)}-F^{*}(t)|\kappa_{n}|^{2}&0\\ 0&0&L_{n}^{(3)}-F(t)|\kappa_{n}|^{2}\end{bmatrix},$
where $L_{n}^{(1)}=\lim_{t\rightarrow\infty}\left(F(t)+F^{*}(t)\right)|\kappa_{n}|^{2}$, $L_{n}^{(2)}=\lim_{t\rightarrow\infty}F^{*}(t)|\kappa_{n}|^{2}$ and $L_{n}^{(3)}=\lim_{t\rightarrow\infty}F(t)|\kappa_{n}|^{2}$, then $\lim_{t\rightarrow\infty}\hat{A}_{n}(t)=\textbf{0}_{3\times 3}$.
For the special case where $g_{n}=0$, meaning the multi-level atom only interacts with the non-Markovian environment with $C_{n}=D_{n}^{T}=\textbf{0}_{1\times 3}$, we then construct the vector representing the populations of a multi-level atom as
$\tilde{X}=\left[\langle\sigma^{+}_{1}\sigma^{-}_{1}\rangle,\cdots,\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle,\cdots,\langle\sigma^{+}_{N-1}\sigma^{-}_{N-1}\rangle\right]^{T},$
which is governed by the following real-value equation if initially $\tilde{X}(0)$ is real,
$\frac{\mathrm{d}}{\mathrm{d}t}\tilde{X}$ $=\begin{bmatrix}-L_{1}^{(1)}+P_{1}&0&\cdots&0\\ 0&-L_{2}^{(1)}+P_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&-L_{N-1}^{(1)}(t)+P_{N-1}\\ \end{bmatrix}\tilde{X}(t)$ (37)
$\triangleq\left(\textbf{L}+\textbf{P}(t)\right)\tilde{X}(t),$
where $\textbf{L}=\mathrm{diag}\left(-L_{1}^{(1)},-L_{2}^{(1)},\cdots,-L_{N-1}^{(1)}\right)$, $\textbf{P}(t)=\mathrm{diag}\left(P_{1},P_{2},\cdots,P_{N-1}\right)$ with $P_{n}=L_{n}^{(1)}-\left(F(t)+F^{*}(t)\right)|\kappa_{n}|^{2}$. The dynamics of Eq. (37) has been analyzed in Ref. [70]. We denote $\tilde{X}(t)=\Phi\left(t,t_{0}\right)\tilde{X}(t_{0})$ for $t\geq t_{0}$, and
$\Phi\left(t,t_{0}\right)=e^{\textbf{L}\left(t-t_{0}\right)}+\int_{t_{0}}^{t}e^{\textbf{L}\left(t-\tau\right)}\textbf{P}(\tau)\Phi\left(\tau,t_{0}\right)\mathrm{d}\tau,$
@@
When the open system converges to a Markvoian regime, the time-varying components in Eq. (37) converge to zero, and the transition matrix is bounded.
@@
The convergence of quantum dynamics to a Markovian regime means the convergence of $F(t)$ according to Definition 1. Consequently, the time-varying components in Eq. (37) converge to zero. In this case, $\lim_{t\rightarrow\infty}\textbf{P}(t)=\textbf{0}_{(N-1)\times(N-1)}$. Thus $\Phi\left(t,t_{0}\right)$ is bounded because $\lim_{t\rightarrow\infty}\frac{\partial\Phi\left(t,t_{0}\right)}{\partial t}=0$. ∎
Obviously, when $\textbf{P}(t)\equiv\textbf{0}_{(N-1)\times(N-1)}$, Eq. (37) reduces to a linear time-invariant system and is exponentially stable because L is a diagonal matrix whose eigenvalues are all negative. Then $\tilde{X}$ can be solved as

### v3

@@ II-B4 Example
Take the parameter settings in Section II-B3) as an example. When $\Omega=\tilde{\omega}_{n}$ and $4\kappa_{n}S_{n}=Q_{n}^{2}$, $I_{n}(t)\equiv 0$, $Q_{n}=-\gamma$, $S_{n}=\frac{\gamma^{2}}{4\kappa_{n}}$, $R_{n}(t)=\frac{\gamma}{2\kappa_{n}}-\frac{\gamma}{\gamma\kappa_{n}t+2\kappa_{n}}$, then $\kappa_{n}R_{n}(t)-\gamma<0$ for arbitrary $t$ and $\textbf{J}\left(X_{F}(t)\right)$ is negative definite. If we take $\mathbf{M}=\mathbf{I}$ as the identity matrix for simplification in Eq. (31), then $\dot{V}(X_{F})<0$ and Eq. (23) is globally stable in this parameter setting.
Generalized from the definition of invariant set as well as bounded-input bounded-output (BIBO) stability defined and introduced in Appendix C, and the relative property in Ref. [51], we can derive the following propositions.
@@
When $\Omega=\tilde{\omega}_{n}$ in Eq. (24) and $\mathbf{M}=\mathbf{I}$ in Eq. (31), there exists $\alpha_{V}>0$ such that the parameters in Eq. (24) satisfy $02\kappa_{n}\alpha_{V}+\gamma\chi_{n}-2\gamma R_{n}<$ 0. In this case, there exists an invariant set of $X_{F}$ represented as $\Omega_{\alpha}=\left\{X_{F}\in\textbf{R}^{2}:V\leq\alpha_{V}\right\}$ contained in the region satisfying $\dot{V}=0$, and $\alpha_{V}\leq\gamma(\gamma-\chi_{n}\kappa_{n})/\kappa_{n}^{2}$.
@@
Assume initially $X_{F}\in\Omega_{\alpha}$ and $\Gamma_{F}$ represents an invariant set of $X_{F}$ satisfying that $\dot{V}(X_{F})=0$. When $02\kappa_{n}\alpha_{V}+\gamma\chi_{n}-2\gamma R_{n}<$ 0, $\dot{V}<0$ is satisfied according to Eqs. (31), (23a) with $\Omega=\tilde{\omega}_{n}$, and $V(x)\geq 0$ by its definition. Then there exists a positive value $v_{l}$ that $\lim_{t\rightarrow\infty}V(X_{F})=v_{l}\geq 0$ and $v_{l}<\alpha_{V}$. Thus $V(X_{F})$ is always within the invariant set. Denote the solutions of $R_{n}$ satisfying $\dot{V}=0$ as $r_{1}$ and $r_{2}$, then $\alpha_{V}\leq r_{1}^{2}+r_{2}^{2}=\gamma(\gamma-\chi_{n}\kappa_{n})/\kappa_{n}^{2}$. ∎
@@
Proposition 1 means that larger $\gamma$ or smaller $\kappa_{n}$ can induce a larger invariant set with stronger Markovian property. This is why the Markovian approximation can be applied in the circumstance that the coupling between the quantum system and the environment is weak.
@@
When $X_{F}$ in Eq. (23) is initially bounded and satisfies that $\dot{V}<0$ when $\Omega=\tilde{\omega}_{n}$, then the nonlinear process in Eq. (23) is BIBO stable.
@@
We take $\mathbf{M}=\mathbf{I}$, $\dot{V}$ is independent of $\Omega-\tilde{\omega}_{n}$. Thus if $\dot{V}<0$ when $\Omega=\tilde{\omega}_{n}$, $\dot{V}<0$ also holds when $\Omega\neq\tilde{\omega}_{n}$ according to Eq. (31), then the system in Eq. (23) is BIBO stable based on Proposition 5 in Appendix C. ∎
@@
When nonlinear process in Eq. (23) is BIBO stable and the Jacobian in Eq. (27) is bounded, the Lyapunov function can also be bounded according to Ref. [52].
@@
For the coupled nonlinear dynamics in Eq. (23), when $\dot{V}<0$ in Eq. (31), $\lim_{t\rightarrow\infty}R_{n}(t)$ and $\lim_{t\rightarrow\infty}I_{n}(t)$ both exist when $\left|\Omega-\tilde{\omega}_{n}\right|\leq\epsilon$.
@@
When $\dot{V}<0$ and $V(t)>0$, $\lim_{t\rightarrow\infty}V(t)$ exists and $\lim_{t\rightarrow\infty}\dot{V}(t)=0$. Then $\lim_{t\rightarrow\infty}R_{n}(t)$ and $\lim_{t\rightarrow\infty}I_{n}(t)$ can be solved by Eq. (31). ∎
Proposition 3.
If the nonlinear dynamics in Eq. (23) is BIBO stable when $\Omega=\tilde{\omega}_{n}$, there exists $\epsilon>0$ such that the trajectory of $X(t)$ in Eq. (9) converges to be Markovian when $t\rightarrow\infty$ if $\left|\Omega-\tilde{\omega}_{n}\right|\leq\epsilon$.
Proof.
Similar to Proposition 2, the Lyapunov function in Eq. (31) is independent of $\left(\Omega-\tilde{\omega}_{n}\right)$. When the system is BIBO stable and $\left|\Omega-\tilde{\omega}_{n}\right|\leq\epsilon$, then $R_{n}(t)$ and $I_{n}(t)$ are solvable with steady values according to Lemma 1 and Eq. (27). Then the trajectory of $X(t)$ in Eq. (9) converges to be Markovian because $R_{n}$ and $I_{n}$ reach constants when $t$ is large enough. ∎
Taking the circumstance $N=2$ as an example, the real and imaginary parts of $F_{1}(t)$ are compared in Fig. 2, where we take $\chi_{1}=1$, $\gamma=2$GHz [53], $\kappa_{1}=1$GHz, $\Omega=50$GHz and $\mathrm{d}t=0.01$ns. The comparisons in Figs. 2(a) and (b) show that, the diverse between $\Omega$ and $\tilde{\omega}_{1}$ can induce oscillations of $R_{1}(t)$ and $I_{1}(t)$, and this can influence the non-Markovian dynamics by via the time-varying decaying rates in Eq. (9).
However, for the multi-level system in Fig. 1, the condition that $\tilde{\omega}_{n}=\Omega$ in Theorem 1 can never be simultaneously satisfied for all the energy levels labeled by $n$, thus we need to generalize Theorem 1 to the circumstance that $\tilde{\omega}_{n}\neq\Omega$ by regarding their difference as uncertain inputs as follows.
II-C Nonlinear dynamics with uncertain input
Practically, we usually cannot know the exact value of the environment parameter $\Omega$, thus the component $\Omega-\tilde{\omega}_{n}$ in Eq. (23) can be regarded as an uncertain input to the nonlinear system. Then Eq. (24) can be re-written as
where we denote $\textbf{u}=\Delta AX_{F}(t)$ with $\Delta A=\begin{bmatrix}0&\Omega-\tilde{\omega}_{n}\\ \tilde{\omega}_{n}-\Omega&0\end{bmatrix}$, and there always exists $\epsilon>0$ satisfying $\|\Delta A\|\leq\epsilon$. Then Eq. (32) can be regarded as a perturbed nonlinear system.
@@ II-C Nonlinear dynamics with uncertain input
The nonlinear equation (23) with $\Omega=\tilde{\omega}_{n}$ is stable according to the parameter settings in Theorem 1.
According to Eqs. (18)-(21), when $\Omega=\tilde{\omega}_{n}$, $I_{n}(t)\equiv 0$, and we denote $\lim_{t\rightarrow\infty}R_{n}(t)=\bar{R}_{n}$. Define a new state vector $\tilde{X}_{F}(t)=\left[R_{n}(t)-\bar{R}_{n},I_{n}(t)\right]^{\rm T}\triangleq\left[\tilde{R}_{n}(t),I_{n}(t)\right]^{\rm T}$ and $\lim_{t\rightarrow\infty}\tilde{R}_{n}(t)=\left[0,0\right]^{\rm T}\triangleq\tilde{X}_{F}^{*}$ when Assumption 1 is satisfied and $\Omega=\tilde{\omega}_{n}$. Then
$\frac{\mathrm{d}}{\mathrm{d}t}\tilde{R}_{n}(t)=\kappa_{n}\left[\left(\tilde{R}_{n}(t)+\bar{I}_{n}\right)^{2}-I_{n}^{2}(t)\right]$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}-\gamma\left(\tilde{R}_{n}(t)+\bar{R}_{n}\right)+\left(\Omega-\tilde{\omega}_{n}\right)I_{n}(t)+\frac{\gamma\chi_{n}}{2},$ (33a)
$\frac{\mathrm{d}}{\mathrm{d}t}I_{n}(t)=2\kappa_{n}\left(\tilde{R}_{n}(t)+\bar{R}_{n}\right)I_{n}(t)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}-\left(\Omega-\tilde{\omega}_{n}\right)\left(\tilde{R}_{n}(t)+\bar{R}_{n}\right)-\gamma I_{n}(t),$ (33b)
and can be simplified as
$\frac{\mathrm{d}}{\mathrm{d}t}\tilde{X}_{F}(t)=$ $\tilde{\textbf{f}}\left(\tilde{X}_{F}(t)\right)+\begin{bmatrix}\frac{\gamma\chi_{n}}{2}+\kappa_{n}\bar{R}_{n}^{2}-\gamma\bar{R}_{n}\\ \left(\tilde{\omega}_{n}-\Omega\right)\bar{R}_{n}\end{bmatrix}$ (34)
$+\begin{bmatrix}0&\Omega-\tilde{\omega}_{n}\\ \tilde{\omega}_{n}-\Omega&0\end{bmatrix}\tilde{X}_{F}(t),$
where $\tilde{\textbf{f}}\left(\tilde{X}_{F}(t)\right)$ is for the nonlinear component on the RHS of Eq. (33). Obviously, $\tilde{\textbf{f}}\left(\tilde{X}_{F}(t)\right)=0$ when $\tilde{X}_{F}=0$. When $\Omega=\tilde{\omega}_{n}$ and $X_{F}(t)$ converges by Theorem 1, the RHS of Eq. (34) converges to zero.
Theorem 2.
When $4\kappa_{n}S_{n}-Q_{n}^{2}\leq 0$ in Eq. (33), there exists $\epsilon>0$ such that the trajectory of $X(t)$ in Eq. (9) converges to a Markovian behavior when the uncertain $\Omega$ is bounded by $\left|\tilde{\omega}_{n}-\Omega\right|\leq\epsilon$.
@@
When $\tilde{\omega}_{n}=\Omega$, by Theorem 1, the quantum dynamics approaches Markovian behavior as $t\rightarrow\infty,$ and the nonlinear dynamics $\frac{\mathrm{d}}{\mathrm{d}t}\tilde{X}_{F}(t)=\tilde{\textbf{f}}\left(\tilde{X}_{F}(t)\right)$ is uniformly and asymptotically stable around its equilibrium $\tilde{X}_{F}=0$. When $\left|\tilde{\omega}_{n}-\Omega\right|\leq\epsilon$ and $\gamma\chi$ is finite, Proposition 6 in Appendix C ensures that $\tilde{R}_{n}(t)$ and $I_{n}(t)$ in Eq. (33) remain stable. Then the trajectory of $X(t)$ in Eq. (9) converges to be Markovian as $t\rightarrow\infty$. ∎
The above theorem shows that, a trajectory of $X(t)$ in Eq. (9) can converge to be Markovian even when $\Omega\neq\tilde{\omega}_{n}$ and $\Omega$ is not precisely known, if only the uncertain $\Omega$ is bounded around $\tilde{\omega}_{n}$.
Now we take $N=3$ as an example for numerical simulations of the above non-Markovian dynamics around the parameter settings in Refs. [54, 55]. As shown in Fig. 3, we take $\omega_{1}=37.7$GHz, $\omega_{2}=75.3$GHz, $\Omega=50$GHz, $\Delta_{1}=-20$MHz, $\Delta_{2}=-70$MHz, $g_{1}=g_{2}=20$MHz, $\chi_{1}=\chi_{2}=1$, $\gamma=10$GHz, $\kappa_{1}=\kappa_{2}=0.31$GHz, and the time step for simulation equals $0.01$ns. In the simulations for Markovian circumstance, we take $F_{n}$ with $n=1,2$ as constants that equal the steady values of those in the non-Markovian simulations.
Figure 3: Non-Markovian and Markovian dynamics of the quantum system with one three-level atom in a cavity.
@@ III Time-varying linear quantum control dynamics
After clarifying the nonlinear parameter dynamics in the section above, in this section, we study the quantum open-loop control and measurement feedback control based on the control model in Eq. (9), and the results are specially influenced by the nonlinear parameter dynamics in Eq. (9e) due to the non-Markovian interactions between the quantum system and environment.
In the following, we only consider the circumstance that $4\kappa_{n}S_{n}-Q_{n}^{2}\leq 0$, thus the interaction between the atom and environment finally converges to a Markovian form. Return to Eq. (12), we analyze how the quantum dynamics can be influenced by detunings between the atom and cavity, and drive fields applied upon the quantum system.
For the simplified circumstance that $\Delta_{n}=0$ as in Refs. [39, 56], then $\tilde{g}_{n}(t)\equiv g_{n}$, $C_{n}$ and $D_{n}$ in Eqs. (15) and (16) are all time-invariant for arbitrary $n$, while $A_{n}$ is time-varying with $F_{n}(t)$ converges to its steady values according to Eqs. (19) and (21). $A(t)$ can be separated into the time-invariant component and time-varying components as
@@
$=$ $\begin{bmatrix}-L_{n}^{(1)}&-i\tilde{g}_{n}&i\tilde{g}_{n}\\ -i\tilde{g}_{n}&-L_{n}^{(2)}&0\\ i\tilde{g}_{n}&0&-L_{n}^{(3)}\end{bmatrix}$
$+{\rm{diag}}\left(L_{n}^{(1)}-\left(F_{n}+F_{n}^{*}\right),L_{n}^{(2)}-F_{n}^{*},L_{n}^{(3)}-F_{n}\right),$
where $L_{n}^{(1)}=\lim_{t\rightarrow\infty}\left(F_{n}(t)+F_{n}^{*}(t)\right)$, $L_{n}^{(2)}=\lim_{t\rightarrow\infty}F_{n}^{*}(t)$ and $L_{n}^{(3)}=\lim_{t\rightarrow\infty}F_{n}(t)$, then $\lim_{t\rightarrow\infty}\hat{A}_{n}(t)=\textbf{0}_{3\times 3}$.
For the special case where $g_{n}=0$, meaning the multi-level atom only interacts with the non-Markovian environment with $C_{n}=D_{n}^{\rm T}=\textbf{0}_{1\times 3}$, we then construct the vector representing the populations of a multi-level atom as
$\tilde{X}=\left[\langle\sigma^{+}_{1}\sigma^{-}_{1}\rangle,\cdots,\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle,\cdots,\langle\sigma^{+}_{N-1}\sigma^{-}_{N-1}\rangle\right]^{\rm T},$
which is governed by the following real-valued equation if initially $\tilde{X}(0)$ is real,
$\frac{\mathrm{d}}{\mathrm{d}t}\tilde{X}=\left(\textbf{L}+\textbf{P}(t)\right)\tilde{X}(t),$ (36)
where $\textbf{P}(n,n)=L_{n}^{(1)}-\left(F_{n}+F_{n}^{*}\right)$, $\textbf{P}(n,n+1)=\left(F_{n+1}+F_{n+1}^{*}\right)-L_{n+1}^{(1)}$, all the other elements in P equal zero, the time-invariant matrix L can be determined by Eq. (35) and $\textbf{P}(t)$, and more details are omitted. We denote $\tilde{X}(t)=\Phi\left(t,t_{0}\right)\tilde{X}(t_{0})$ for $t\geq t_{0}$, and [57]
$\Phi\left(t,t_{0}\right)=e^{\textbf{L}\left(t-t_{0}\right)}+\int_{t_{0}}^{t}e^{\textbf{L}\left(t-\tau\right)}\textbf{P}(\tau)\Phi\left(\tau,t_{0}\right)\mathrm{d}\tau,$ (37)
@@
When the trajectories of $\tilde{X}(t)$ in Eq. (36) with $g_{n}=0$ converge to be Markovian, the time-varying component P converges to zero, and the transition matrix in Eq. (37) is bounded.
@@
The convergence of quantum dynamics to a Markovian regime means the convergence of $F_{n}(t)$ according to Remark 1. Consequently, the time-varying components in Eq. (36) converge to zero. In this case, $\lim_{t\rightarrow\infty}\textbf{P}(t)=\textbf{0}_{(N-1)\times(N-1)}$. Thus $\Phi\left(t,t_{0}\right)$ is bounded because $\lim_{t\rightarrow\infty}\frac{\partial\Phi\left(t,t_{0}\right)}{\partial t}=0$. ∎
Specially, when $\textbf{P}(t)\equiv\textbf{0}_{(N-1)\times(N-1)}$, Eq. (36) reduces to a linear time-invariant system and is exponentially stable because L is a diagonal matrix whose eigenvalues are all negative.
Generally, $\tilde{X}$ can be solved as